# LEMMA-RCA: A Large Multi-modal Multi-domain Dataset for Root Cause Analysis

### v1

@@ Abstract
Root cause analysis (RCA) is crucial for enhancing the reliability and performance of complex systems. However, progress in this field has been hindered by the lack of large-scale, open-source datasets tailored for RCA. To bridge this gap, we introduce LEMMA-RCA, a large dataset designed for diverse RCA tasks across multiple domains and modalities. LEMMA-RCA features various real-world fault scenarios from IT and OT operation systems, encompassing microservices, water distribution, and water treatment systems, with hundreds of system entities involved. We evaluate the quality of LEMMA-RCA by testing the performance of eight baseline methods on this dataset under various settings, including offline and online modes as well as single and multiple modalities. Our experimental results demonstrate the high quality of LEMMA-RCA. The dataset is publicly available at https://lemma-rca.github.io/.
@@ 1 Introduction
Root cause analysis has been extensively studied across various domains and settings [1; 2; 3; 4; 5]. Based on the application scenarios, RCA can be carried out in offline/online fashion with single/multi-modal system data. Existing studies on RCA in these settings involve numerous learning techniques such as Bayesian methods [6], decision trees [7], etc. Particularly, causal structure learning technique [8; 9; 10; 11; 12; 13; 14; 15] has proven effective in constructing causal or dependency graphs between different system entities and key performance indicators (KPIs), thereby enabling the tracing of underlying causes through these structures.
@@
To address these limitations, we introduce LEMMA-RCA, a collection of Large-scalE Multi-ModAl datasets with various real system faults to facilitate future research in Root Cause Analysis. LEMMA-RCA is multi-domain, encompassing real-world applications such as IT operations and water treatment systems, with hundreds of system entities involved. LEMMA-RCA accommodates multi-modal data including textual system logs with millions of event records and time series metric data with more than $100,000$ timestamps. We annotate LEMMA-RCA with ground truth labels indicating the precise time stamps when real system faults occur and their corresponding root-cause system entities.
@@
 Dataset Public Real Faults Large-scale Multi-domain Modality Single Multiple LEMMA-RCA NeZha PetShop ITOps Murphy
A comparison between LEMMA-RCA and existing datasets for RCA is presented in Table 1. We briefly discuss the status of existing datasets: 1) NeZha [12] has limited size and contains many missing parts in the monitoring data, and it is confined to one domain: microservice architectures. 2) PetShop [17] has a small size. Additionally, the system comprises only $41$ components, limiting its complexity and reducing the practicality for real-world scenarios. 3) ITOps [18] dataset is not public and contains structured logs that do not contribute to comprehending the underlying causal mechanism of system failures, making it difficult to conduct fine-grained RCA. 4) Murphy [16] is collected from a simple system and also not public. In comparison to prior work, LEMMA-RCA demonstrates a comprehensive maturity on the accessibility, authenticity, and diversity.
LEMMA-RCA enables fair comparisons among different RCA methods. We evaluate eight baseline methods, with five suited for offline settings and the remaining three designed for online RCA. The quality of various data modalities is assessed in both online and offline setups. As previously mentioned, causal-graph-based RCA methods provide deeper insights into system failures; hence, all baseline methods fall into this category. The experimental results demonstrate the high quality of LEMMA-RCA and its extensive utility for advanced research in root cause analysis.
2 Preliminaries
@@
Data-driven Root Cause Analysis Problem. Given the monitoring data (including metrics and logs) of system entities and system KPIs, the root cause analysis problem is to identify the top $K$ system entities that are most relevant to KPIs when a system fault occurs. RCA techniques can be implemented in various settings, where offline/online and single-modal/multi-modal are mostly commonly concerned. Offline RCA is conducted retrospectively with historical data to determine past failures, whereas online RCA operates in real-time using current data streams to promptly address issues. On the other hand, single-modal RCA relies solely on one type of data for a focused analysis, while multi-modal RCA uses multiple data sources for a comprehensive assessment. We illustrate the procedure of RCA in single-modal offline and multi-modal online settings in Figure 1.
Figure 1: Illustration of RCA workflow in the single-modal offline setting (top) and the multi-modal online setting (bottom). The other two settings can be viewed as an ensemble of corresponding components (data collection, detector, modality) and follow the same systematic procedure.
3 LEMMA-RCA Dataset
This section outlines the data resources, details the preprocessing steps, and presents visualizations to illustrate the characteristics of the data released.
@@ 3.1 Data Collection
We collect real-world data from two domains: IT operations and OT operations. The IT domain includes sub-datasets from Product Review and Cloud Computing microservice systems, while the OT domain includes SWaT and WADI sub-datasets from water treatment and distribution systems. Data specifics are provided in Table 2 and Table 3.
@@
Figure 2: Visualization of the microservice system platform and ElasticSearch log data.
@@
In the IT domain, we developed two microservice platforms: the Product Review Platform and the Cloud Computing Platform. The Product Review Platform is composed of six OpenShift nodes (such as ocp4-control-plane-1 through ocp4-control-plane-3, ocp4-compute-1 and ocp4-compute-2, and ocp4-infra-1) and $216$ system pods (including ProductPage, MongoDB, review, rating, payment, Catalogue, shipping, etc.). In this setup, we simulated four distinct system faults, including out-of-memory, high-CPU-usage, external-storage-full, and DDoS attack, on four different dates. Each simulation operated the microservice system for at least $49$ hours with different pods involved. The structure of this microservice system with some key pods of one simulation is depicted Figure 2 (a). Both log and metric data were generated and stored systematically to ensure comprehensive monitoring. Specifically, eleven types of node-level metrics (e.g., net disk IO usage, net disk space usage, etc.) and six types of pod-level metrics (e.g., CPU usage, memory usage, etc.) were recorded by Prometheus [19], and the time granularity of these system metrics is 1 second. Log data, on the other hand, were collected by ElasticSearch [20] and stored in JSON files with detailed timestamps and retrieval periods. The contents of system logs include timestamp, pod name, log message, etc., as shown in Figure 2 (b). The JMeter [21] was employed to collect the system status information, such as elapsed time, latency, connect time, thread name, throughput, etc. The latency is considered as system KPI as the system failure would result in the latency significantly increasing.
@@
For the Cloud Computing Platform, we simulated six different types of faults (such as cryptojacking, mistakes made by GitOps, configuration change failure, etc.) on eleven system nodes. At each simulation, we gathered both system metrics and logs from various sources. In contrast to the Product Review platform, system metrics were directly extracted from CloudWatch Metrics on EC2 instances, and the time granularity of these system metrics is 1 second. Log events were acquired from CloudWatch Logs, consisting of three data types (i.e., log messages, api debug log, and mysql log). Log message describes general log message about all system entities; api debug log contains debug information of the AP layer when the API was executed; mysql logs contain log information from database layer, including connection logs to mysql, which user connected from which host, and what queries were executed. Latency, error rate, and utilization rate were tracked using JMeter tool, serving as Key performance indicators (KPIs). This comprehensive logging and data storage setup facilitates detailed monitoring and analysis of the system’s performance and behavior.
Table 2: Data statistics of IT operation sub-datasets.
@@
 Microservice System Product Review Cloud Computing Original Dataset Size 765 GB 540 GB Number of fault types 4 6 Average number of entities per fault 216.0 167.71 Average number of metrics per fault 11 (node-level) + 6 (pod-level) 6 (node-level) + 7 (pod-level) Average number of timestamps per fault 131,329.25 109,350.57 Average max log events per fault across pods 153,081,219.0 63,768,587.25
@@
Table 3: Data statistics of OT operation sub-datasets.
 Water Treatment/Distribution SWaT WADI Original Dataset Size 235.5 MB 847.6 MB Number of fault types 16 9 Average number of entities per fault 51.0 123.0 Average number of metrics per fault 7 (node-level) + 7 (pod-level) 7 (node-level) + 7 (pod-level) Average number of timestamps per fault 56239.88 85248.47
@@ 3.2 Data Preprocessing
KPI Construction for SWaT and WADI. The SWaT and WADI sub-datasets include the label column that reflects the system status; however, the values within this column are discrete. To facilitate the root cause analysis, it is beneficial to transform these values into a continuous format. Specifically, we propose to convert the label into a continuous time series. To achieve this, we employ anomaly detection algorithms, such as Support Vector Data Description and Isolation Forest, to model the data. Subsequently, the anomaly score, as determined by the model, will be utilized as the system KPI.
@@ 3.3 System Fault Scenarios
We simulate $10$ different types of real system faults in Product Review and Cloud Computing platforms. Due to the space limitation, we select two representative cases (one from each) and provide the details below. We also visualize the system fault of these two cases in Figure 4.
@@
  • External Storage Failure. In this system failure, we fill up the external storage disk connected to the Database (DB) pod (i.e., mongodb-v1) within Microservice A’s OpenShift cluster. When the storage becomes full, the DB pod cannot add new data, resulting in system errors. These errors propagate to pods that depend on the DB pod, causing some services (ratings) within Microservice A to encounter errors. We monitor changes in response and error information for Microservice A using Jaeger logs. Metrics for all containers and nodes, including CPU and memory usage, are obtained from Prometheus within OpenShift. Logs for all containers and nodes are retrieved from Elasticsearch within OpenShift. Additionally, we collect message logs from the external storage. We illustrate the metrics and log data of the root cause pod in Figure 5.
@@ 4.1 Experimental Setup
Evaluation Metrics. To asses baseline RCA method on LEMMA-RCA, we choose three widely-used metrics [14; 25; 15] and introduce them below.
@@
where $rank_{R_{a}}$ is the rank number of the first correctly predicted root cause for system fault $a$.
Baselines. We evaluate the performance of the following causal discovery based RCA models on the benchmark sub-datasets: (1). PC [8]: This classic constraint-based causal discovery algorithm is designed to identify the causal graph’s skeleton using an independence test. (2) Dynotears [9]: It constructs dynamic Bayesian networks through vector autoregression models. (3). C-LSTM [11]: This model utilizes LSTM to model temporal dependencies and capture nonlinear Granger causality. (4). GOLEM [10]: GOLEM relaxes the hard Directed Acyclic Graph (DAG) constraint of NOTEARS [26] with a scoring function. (5). REASON [14]: An interdependent network model learning both intra-level and inter-level causal relationships. (6). Nezha [12]: A multi-modal method designed to identify root causes by detecting abnormal patterns. (7). MULAN [15]: A multi-modal RCA method that learns the correlation between different modalities and co-constructs a causal graph for root cause identification. (8). CORAL [13]: An online single-modal RCA method based on incremental disentangled causal graph learning.
The fist four models can only learn the causal structure from time series data. Thus, we first collect monitoring data from the beginning until system failures occur as historical records. Then, based on the collected records, we apply the causal discovery models to learn causal graphs and leverage random walk with restarts on such graphs [13] to identify the top $K$ nodes as the root causes. Besides we extend NOTEARS and GOLEM to the online learning setting, denoted by NOTEARS and GOLEM, respectively [^1] Other baselines are not extended to the online setting as they are time-intensive when there are multiple data batches.. For the online setting, we use the historical normal data (e.g., $8$ hours for the Product Review sub-dataset, and $1$ hour for the SWaT and WADI sub-datasets) to construct the initial causal graph and update iteratively for each new batch of data. CORAL can inherit the causations from the previous data batch, while NOTEARS and GOLEM have to learn from scratch for each new data batch.
4.2 Offline Root Cause Analysis Results
Product Review and Cloud Computing. We evaluate seven offline RCA methods including both single-modal and multi-modal methods on Product Review and Cloud Computing sub-dataset. The experimental results are presented in Table 4 with respect to Precision at K (PR@K), Mean Reciprocal Rank (MRR), and Mean Average Precision at K (MAP@K). Our observations reveal the following insights: (1) The REASON method demonstrates notable success in identifying the root cause in 75% of system fault scenarios, achieving a PR@1 score of 75%. This indicates the utility of metric data alone in facilitating root cause identification. (2) The performance of these RCA methods is diminished when relying solely on log data for root cause analysis. (3) Integrating both metric and log data enhances the performance of most RCA methods in terms of MRR, compared to using only metric data. This suggests that log data complements these methods, aiding in more accurate identification of potential root causes.
@@ 4.2 Offline Root Cause Analysis Results
Table 4: Results for offline RCA baselines with multiple modalities on the Product Review dataset.

### v3

@@ Abstract
Root cause analysis (RCA) is crucial for enhancing the reliability and performance of complex systems. However, progress in this field has been hindered by the lack of large-scale, open-source datasets tailored for RCA. To bridge this gap, we introduce LEMMA-RCA, a large dataset designed for diverse RCA tasks across multiple domains and modalities. LEMMA-RCA features various real-world fault scenarios from Information Technology (IT) and Operational Technology (OT) systems, encompassing microservices, water distribution, and water treatment systems, with hundreds of system entities involved. We evaluate the performance of six baseline methods on LEMMA-RCA under various single- and multi-modal configurations. Our study demonstrates the utility of LEMMA-RCA in facilitating fair evaluation and promoting the development of more robust RCA techniques. The dataset and code are publicly available at https://lemma-rca.github.io/.
@@ 1 Introduction
Root cause analysis has been extensively studied across various domains and settings [1, 2, 3, 4, 5]. Depending on the nature of the system and the available data, RCA can utilize either single-modal or multi-modal information sources, encompassing logs, metrics, or domain-specific indicators. A diverse array of analytical and machine learning techniques have been employed for RCA, reflecting the complexity and variability of real-world fault diagnosis tasks. These techniques include probabilistic models such as Bayesian networks [6], interpretable methods like decision trees [7], and graph-based reasoning approaches [8, 9, 10] that capture interdependencies among system components.
@@
To address these limitations, we introduce LEMMA-RCA, a collection of Large-scalE Multi-ModAl datasets with various real system faults to facilitate future research in Root Cause Analysis. LEMMA-RCA encompasses real-world applications such as IT operations and water treatment systems, with hundreds of system entities involved. LEMMA-RCA accommodates multi-modal data including textual system logs with millions of event records and time series metric data with more than $100,000$ timestamps. We annotate LEMMA-RCA with ground truth labels indicating the precise time stamps when real system faults occur and their corresponding root-cause system entities.
@@
 Dataset Public Real Faults Large-scale Multi-domain Modality Single Multiple LEMMA-RCA NeZha PetShop Sock-Shop ITOps Murphy
A comparison between LEMMA-RCA and existing datasets for RCA is presented in Table 1. We briefly discuss the status of existing datasets: 1) NeZha [10] has limited size and contains many missing parts in the monitoring data, and it is confined to one domain: microservice architectures. 2) PetShop [12] has a small size. Additionally, the system comprises only $41$ components, limiting its complexity and reducing the practicality for real-world scenarios. 3) Sock-Shop [13] is small-scale with only 13 microservices, and the injected faults (CPU hog and memory leak) are synthetic. Additionally, the data is not publicly available and consists solely of single-modality metrics, lacking diversity in data sources such as logs or traces. 4) ITOps [14] dataset is not public and contains structured logs that offer limited insight into the underlying failure mechanisms, making it challenging to perform fine-grained root cause analysis.5) Murphy [11] is collected from a simple system and also not public. In comparison to prior work, LEMMA-RCA demonstrates a comprehensive maturity on the accessibility, authenticity, and diversity.
LEMMA-RCA enables fair and comprehensive comparisons across a variety of RCA methods. We evaluate six baseline algorithms and assess the quality of different data modalities. The experimental results demonstrate the effectiveness of LEMMA-RCA in benchmarking RCA methods and highlight its broad potential to support advanced research in root cause analysis.
2 Preliminaries and Related Work
@@
Data-driven Root Cause Analysis Problem. Given the monitoring data (including metrics and logs) of system entities and system KPIs, the root cause analysis problem is to identify the top $K$ system entities that are most relevant to KPIs when a system fault occurs. RCA techniques can be implemented in various settings, where single-modal/multi-modal are mostly commonly concerned. Single-modal RCA relies solely on one type of data for a focused analysis, while multi-modal RCA uses multiple data sources for a comprehensive assessment.
Single-modal Root Cause Analysis (RCA) retrospectively identifies the primary cause of system failures using a single data type after an event has occurred [15, 16, 17, 18, 19]. For example, Meng et al. [17] analyze monitoring metric data to discern sequential relationships and integrate causal and temporal information for root cause localization in microservice systems. Similarly, Wang et al. [15] construct an interdependent causal network from time series data, using a random walk strategy to pinpoint the most probable root causes. Li et al. [18] evaluate microservice traces, determining that a service with a higher ratio of abnormal to normal traces is likely the root cause. Recently, large language model (LLM) based methods become a new research direction to learn causal relation for root cause identification due to the success of LLMs in performing complex tasks [20, 21, 22, 23, 24, 25]. For instance, Chen et al. [20] introduce RCACopilot, an innovative on-call system empowered by the large language model for automating RCA of cloud incidents. Shan et al. [21] propose to first identify the log messages indicating configuration-related errors and then localize the suspected root-cause configuration properties based on the selected log messages and the offered configuration settings by LLMs. Goel et al. [22] demonstrate that LLMs can benefit from service functionality and upstream dependency information in better reasoning, thus improving the quality of the identification of root causes. Although these studies demonstrate notable efficacy, they rely exclusively on single-modal data, which may lead to suboptimal and biased outcomes in root cause localization.
Multi-modal RCA. Recent studies have explored utilizing multi-modal data for RCA, which can be divided into two approaches [10, 26, 27, 28]. The first approach, exemplified by Nezha [10] and PDiagnose [26], involves extracting information from each modality separately and then integrating it for analysis. Conversely, the second approach focuses on the interactions between different modalities. For instance, MULAN [27] develops a comprehensive dependency graph by learning correlations between modalities, while MM-DAG [28] aims to jointly learn multiple Direct Acyclic Graphs, improving both consistency and depth of analysis.
3 LEMMA-RCA Data
This section outlines the data resources, details the preprocessing steps, and presents visualizations to illustrate the characteristics of the data released. The data licence can be found in appendix C.
@@ 3.1 Data Collection
We collect real-world data from two domains: IT operations and OT operations. The IT domain includes sub-datasets from Product Review and Cloud Computing microservice systems, while the OT domain includes Secure Water Treatment (SWaT) and Water Distribution (WADI) sub-datasets from water treatment and distribution systems. Data specifics are provided in Table 2.
@@
Figure 1: Visualization of the microservice system platform, which contains 6 nodes and multiple pods that may vary across different stages; and the ElasticSearch log data.
@@
In the IT domain, we developed two microservice platforms: the Product Review Platform and the Cloud Computing Platform. The Product Review Platform is composed of six OpenShift nodes (such as ocp4-control-plane-1 through ocp4-control-plane-3, ocp4-compute-1 and ocp4-compute-2, and ocp4-infra-1) and $216$ system pods (including ProductPage, MongoDB, review, rating, payment, Catalogue, shipping, etc.). In this setup, four distinct system faults are colleceted, including out-of-memory, high-CPU-usage, external-storage-full, and DDoS attack, on four different dates. Each system fault ran the microservice system for at least $49$ hours with different pods involved. The pods running in different stages may vary, and the pods associated with different types of system faults also differ. The structure of this microservice system with some key pods during one fault is depicted in Figure 1 (a). Both log and metric data were generated and stored systematically to ensure comprehensive monitoring. Specifically, eleven types of node-level metrics (e.g., net disk IO usage, net disk space usage, etc.) and six types of pod-level metrics (e.g., CPU usage, memory usage, etc.) were recorded by Prometheus [29], and the time granularity of these system metrics is 1 second. Log data, on the other hand, were collected by ElasticSearch [30] and stored in JSON files with detailed timestamps and retrieval periods. The contents of system logs include timestamp, pod name, log message, etc., as shown in Figure 1 (b). The JMeter [31] was employed to collect the system status information, such as elapsed time, latency, connect time, thread name, throughput, etc. The latency is considered as system KPI as the system failure would result in the latency significantly increasing.
@@
For the Cloud Computing Platform, we monitored six different types of faults (such as cryptojacking, mistakes made by GitOps, configuration change failure, etc.), and collected system metrics and logs from various sources. In contrast to the Product Review platform, system metrics were directly extracted from CloudWatch [^1] https://aws.amazon.com/cloudwatch/ Metrics on EC2 instances, and the time granularity of these system metrics is 1 second. Log events were acquired from CloudWatch Logs, consisting of three data types (i.e., log messages, api debug log, and mysql log). Log message describes general log message about all system entities; api debug log contains debug information of the AP layer when the API was executed; mysql logs contain log information from database layer, including connection logs to mysql, which user connected from which host, and what queries were executed. Latency, error rate, and utilization rate were tracked using JMeter tool, serving as Key performance indicators (KPIs). This comprehensive logging and data storage setup facilitates detailed monitoring and analysis of the system’s performance and behavior.
Table 2: Data statistics of IT and OT operation sub-datasets.
@@
 Microservice System (IT) Product Review Cloud Computing Original Dataset Size 765 GB 540 GB Number of (#) fault types 4 6 Average # entities per fault 216.0 167.71 Average # metrics per fault 11 (node-level) + 6 (pod-level) 6 (node-level) + 7 (pod-level) Average # timestamps per fault 131,329.25 109,350.57 Average max log events per fault across pods 153,081,219.0 63,768,587.25  Water Treatment/Distribution (OT) SWaT WADI Original Dataset Size 4.47G 5.67G Number of (#)fault types 16 9 Average # entities per fault 51.0 123.0 Average # metrics per fault 7 (node-level) + 7 (pod-level) 7 (node-level) + 7 (pod-level) Average # timestamps per fault 56239.88 85248.47
@@ 3.2 Data Preprocessing
KPI Construction for SWaT and WADI. The SWaT and WADI sub-datasets include the label column that reflects the system status; however, the values within this column are discrete. To facilitate the root cause analysis, it is beneficial to transform these values into a continuous format. Specifically, we propose to convert the label into a continuous time series. To achieve this, we employ anomaly detection algorithms, such as Support Vector Data Description and Isolation Forest, to model the data. Subsequently, the anomaly score, as determined by the model, will be utilized as the system KPI. More data preprocessing details on SWaT and WADI can be found in Appendix A
@@ 3.3 System Fault Scenarios
There are $10$ different types of real system faults in Product Review and Cloud Computing sub-datasets. Due to the space limitation, we select two representative cases (one from each) and provide the details below. Other fault scenarios are presented in Appendix B. We also visualize the system fault of these two cases in Figure 3.
@@
  • External Storage Failure. In this system failure, we fill up the external storage disk connected to the Database (DB) pod (i.e., mongodb-v1) within Microservice A’s OpenShift [^2] https://www.redhat.com/en/technologies/cloud-computing/openshift cluster. When the storage becomes full, the DB pod cannot add new data, resulting in system errors. These errors propagate to pods that depend on the DB pod, causing some services (ratings) within Microservice A to encounter errors. We monitor changes in response and error information for Microservice A using Jaeger logs. Metrics for all containers and nodes, including CPU and memory usage, are obtained from Prometheus within OpenShift. Logs for all containers and nodes are retrieved from Elasticsearch within OpenShift. Additionally, we collect message logs from the external storage. We illustrate the metrics and log data of the root cause pod in Figure 4.
@@ 4.1 Experimental Setup
Evaluation Metrics. To assess baseline RCA method on LEMMA-RCA, we choose three widely-used metrics [15, 9, 27] and introduce them below.
@@
where $\text{rank}_{R_{a}}$ is the rank number of the first correctly predicted root cause for system fault $a$.
Baselines. We evaluate the performance of the following RCA baselines on the benchmark sub-datasets, selecting only those with publicly available code to ensure fair and reproducible comparisons: (1). PC-based [35]: This approach first employs a Peter-Clark (PC) algorithm [36] to construct the anomaly behavior graph, and then applies a random walk algorithm to rank the root causes based on the estimated graph structure. (2). CIRCA [37]: This model utilizes structural graph construction, regression-based hypothesis testing, and descendant adjustment to identify root cause metrics. (3). $\epsilon$-Diagnosis [38]: This model diagnoses small-window, long-tail latency in large-scale microservice platforms using a two-sample test and $\epsilon$-statistics. (4). RCD [13]: This technique hierarchically localizes the root cause of failures by focusing on relevant sections of the dependency graph. (5). BARO [39]: It is an end-to-end approach integrating Bayesian change point detection and nonparametric hypothesis testing to accurately detect anomalies and identify root causes in microservice systems. (6). Nezha [10]: A multi-modal method designed to identify root causes by detecting abnormal patterns.
When using only metric data (e.g., CPU usage, memory usage) for root cause analysis, we first identify the potential root cause associated with each metric. We then compute the final score for each candidate by averaging its ranking scores across all metrics. This procedure is similarly applied in log-only and multi-modal scenarios to rank and identify the root causes. For the hyperparameters, we use the default parameter values for all baselines to ensure a fair comparison.
@@
Table 3: Results for RCA baselines with multiple modalities on the Product Review dataset.